OpenStudy (anonymous):

I = 4000 x ( 9 + 1)/100 x 3

OpenStudy (anonymous):

I= 120

OpenStudy (mathstudent55):

Is this the problem? \(I = 4000 \times \dfrac{9 + 1}{100} \times 3\)

OpenStudy (mathstudent55):

This is what you wrote above, so I'll help you with it. \(I = 4000 \times \dfrac{9 + 1}{100} \times 3\) First, you must do the addition in the numerator. \(I = 4000 \times \dfrac{10}{100} \times 3\) Now do the division in the fraction. \(I = 4000 \times 0.1 \times 3\) Now multiply the three numbers together. \(I = 1200\)